OpenStudy (anonymous):

Find the minimum and maximum for P = 3x + y. Vertices are (0,0), (4, 0), (3, 4), (0, 10). Show all work.

OpenStudy (anonymous):

@oldrin.bataku ? :b

OpenStudy (anonymous):

\[ P(0,0)=3(0)+0=0\\ P(4,0)=3(4)+0=12\\ P(3,4)=3(3)+4=13\\ P(0,10)=3(0)+10=10 \] The minimum is at (0,0) while the maximum is at (3,4)

OpenStudy (anonymous):

But how did you figure that out? I don't want just the answer, I need to see how you did it

OpenStudy (anonymous):

@oldrin.bataku ? How did you figure it out??